# Benchmarking multi-step methods for the dynamic prediction of survival with numerous longitudinal predictors

### v2

@@ 1.1 Dynamic prediction
1.2 Traditional dynamic prediction methods
Traditional approaches to the dynamic prediction of survival outcomes include landmarking (Van Houwelingen, , 2007) and joint modelling (Hickey et al., , 2016).
The landmarking method specifies a conditional model for the probability of experiencing the event in the future, given that an individual has not yet experienced the event up until a prespecified landmark time. Individuals that either experienced the event or were censored before the landmark are discarded, while individuals who are still at risk at the landmark time are retained for analysis. Then, all repeated measurements of each longitudinal covariate taken up until the landmark time are summarized into a single summary measure, and the summary measures of the longitudinal covariates are used as predictors in a Cox proportional hazards model (Cox, , 1972) alongside the baseline covariates. In the simplest and most commonly used approach, called last observation carried forward (LOCF) landmarking, the last observation of each longitudinal covariate before the landmark time is used as summary measure. The major advantages of the LOFC approach are its simplicity and ease of implementation; possible disadvantages include the fact that the last observation taken before the landmark may be outdated, and the lack of a model to describe the evolution over time of the longitudinal covariates and to account for measurement error (which may be considerable in the case of biomarkers).
@@ 1.2 Traditional dynamic prediction methods
Joint modelling involves the specification and estimation of a shared random effects model (called “joint model”) for both the longitudinal covariates and the survival outcomes. Usually, the joint model (JM) assumes the survival outcome and the longitudinal covariates to be conditionally independent given a vector of subject-specific random effects. Unlike LOCF landmarking, JMs enable efficient use of the longitudinal data, and to account for measurement error. However, their estimation is computationally difficult and demanding, particularly when attempting to include many longitudinal covariates in the joint model. To date, high computing times and frequent convergence problems have limited the applicability of JMs to problems with a limited number of longitudinal covariates (Fournier et al., , 2019; Spreafico and Ieva, , 2021), often forcing researchers working on data with numerous longitudinal covariates to preselect a subset of those covariates before fitting the JM.
1.3 Multi-step dynamic prediction methods
Over the last decades, technological advancements have led to a growing availability of studies where many longitudinal covariates are measured alongside a survival outcome. Examples include observational studies were patients are monitored for several years to study the occurrence of a slowly progressing disease such as cancer (Patel et al., , 2017), dementia (Bennett et al., , 2018) or Alzheimer’s disease (Weiner et al., , 2010), as well as studies that seek to identify biomarkers associated with disease progression and the occurrence of disease milestones among a large number of longitudinal omic variables (Signorelli et al., , 2020; Filbin et al., , 2021).
@@ 1.3 Multi-step dynamic prediction methods
Recently, this increasing availability of repeated measurement data in biomedical studies has driven the development of new statistical methods (Li and Luo, , 2019; Signorelli et al., , 2021; Lin et al., , 2021; Devaux et al., , 2023; Gomon et al., , 2024) and software (Signorelli, , 2024; Devaux et al., , 2024) for dynamic prediction that can better handle the availability of a large number of longitudinal covariates. These methods approach the dynamic prediction problem through a multi-step modelling approach: first, they employ either mixed-effects models (MEMs) or multivariate functional principal component analysis (MFPCA) to describe the evolution over time of the longitudinal covariates and to obtain time-independent summaries of the repeated measurement data; then, they model the relationship between the survival outcome and the available covariates using either Cox models or random survival forests (RSFs).
In particular, the Multivariate Functional Principal Component Analysis Cox model (MFPCCox, Li and Luo, (2019); Gomon et al., (2024)) uses MFPCA to reduce the longitudinal covariates into subject-specific MFPCA scores, and then proceeds to estimate a Cox model where both baseline covariates and the MFPCA scores are included as covariates. Penalized Regression Calibration (PRC, Signorelli et al., (2021); Signorelli, (2024)) makes use of linear mixed models (LMMs) as well as multivariate mixed models to reduce the longitudinal covariates into subject-specific random effects, then it employs a penalized Cox model to predict the survival outcome based on the predicted random effects and baseline covariates. Similarly to MFPCCox, the Functional Random Survival Forest (FunRSF, Lin et al., (2021)) uses MFPCA to obtain MFPCA scores, but then it uses a random survival forest to predict survival based on the MFPCA scores and baseline covariates. Lastly, similarly to PRC, the dynamic Random Survival Forest (DynForest, Devaux et al., (2023, 2024)) method uses LMMs to compute the predicted random effects; however, it uses the predicted random effects and baseline covariates to predict survival through a random survival forest.
1.4 Benchmarking multi-step dynamic prediction methods
@@
The goal of this study is to investigate these open questions using publicly available real-world datasets. We gathered data from three longitudinal studies that differ substantially in terms of sample size, number of longitudinal covariates and length of follow-up. We used MFPCCox, PRC, FunSRF and DynForest, alongside two simpler prediction approaches, to predict a relevant time-to-event outcome in each dataset, and proceeded to evaluate the predictive performance of each method according to three different metrics for survival data. Although an additional comparison with the JM approach would be interesting, JMs are not included in this benchmarking study because their estimation through state of the art statistical software for JMs (such as joineMRL and JMbayes2) repeatedly failed when attempting to include more than three to five longitudinal predictors in the JM.
@@ 2.1 Dynamic prediction of survival outcomes
We consider the setup of a longitudinal study where $n$ subjects are enrolled, and interest lies in predicting a time-to-event outcome. Each subject $i\in\{1,2,...,n\}$ is followed from time $t=0$ (study entry / baseline visit) until they either experience the event of interest at time $t_{i}$, or they are right-censored at time $c_{i}$. We denote by $T_{i}$ and $C_{i}$ the random variables corresponding to the event and censoring times, respectively, and let $T_{i}^{*}=\min{\left(T_{i},C_{i}\right)}$. Furthermore, we denote by $t_{i}^{*}$ the observed value of $T_{i}^{*}$, and introduce the censoring indicator $\delta_{i}$ that is $1$ if for subject $i$ the event of interest is observed at $t_{i}^{*}=t_{i}$, and 00 in case of censoring, i.e., if the event has not yet been observed at $t_{i}^{*}=c_{i}$.
@@ 2.3.6 Dynamic Random Survival Forest (DynForest)
2.3.7 Statistical software
The static Cox model and LOCF landmarking model can be estimated using the coxph function from R package survival (Therneau and Grambsch, , 2000). Dedicated R packages are available for PRC and DynForest: PRC can be estimated using the R package pencal (Signorelli, , 2024), and DynForest using the R package DynForest (Devaux et al., , 2024). As concerns the remaining two methods, MFPCCox can be estimated using the R code provided in Li and Luo, (2019) and FunRSF using the R code provided in Lin et al., (2021).
All computations presented in this study were performed on an AMD EPYC 7662 processor with 2 GHz CPU.
@@ 3.4.2 Performance measures and validation of predictive performance
We employed the Brier score (Graf et al., , 1999), time-dependent AUC (Heagerty et al., , 2000), and C index (Pencina and D’Agostino, , 2004) as measures of predictive performance. We employed repeated cross-validation (RCV) to obtain unbiased estimates of the predictive performance of the different models. For the ROSMAP and ADNI datasets we employed 10-fold cross-validation (CV) with 10 repetitions; for the PBC2 dataset, where using 10-fold CV led to very small validation folds, we employed 5-fold CV with 20 repetitions.
@@ 3.4.3 Implementation details
[…]
@@ 4.4 Computing time
We now turn our attention to computing time. Tables 4, 5, and 6 report the average computing time required to estimate each model in $k-1$ folds and to obtain predictions for the remaining fold (averaged over all RCV iterations). Computing time is generally higher for the ROSMAP dataset, where we have the largest sample size and number of predictors, and lowest for the PBC2 dataset where $n$, $P$ and $Q$ are small. By comparing computing time across landmarks, we observe that the computing time typically decreases as the landmark increases; this is simply an artifact of the fact that at later landmarks less individuals are still at risk, reducing the sample size used for model training and prediction.
@@ 5.3 Software and computing time
Supplementary material and R code
Supplementary tables and figures are available in the “Supplementary Material” pdf file. The R code used to perform the analyses described in this paper is available at
https://github.com/mirkosignorelli/comparisonDynamicPred.
@@ Acknowledgements
ROSMAP. ROSMAP is supported by P30AG10161, P30AG72975, R01AG15819, R01AG17917, U01AG46152, and U01AG61356.
ADNI. Data from the ADNI study were obtained from the ADNI database (http://adni.loni.usc.edu). As such, the investigators within the ADNI contributed to the design and implementation of ADNI and/or provided data but did not participate in analysis or writing of this paper. A complete listing of ADNI investigators can be found at https://adni.loni.usc.edu/wp-content/uploads/how_to_apply/ADNI_Acknowledgement_List.pdf. Data from the ROSMAP study were obtained from the Rush Alzheimer’s Disease Center (RADC) of the RUSH University (https://www.radc.rush.edu).

### v3

@@ Abstract
About this article
  • Please cite this article as: Signorelli, M., Retif, S. (2025). Benchmarking multi-step methods for the dynamic prediction of survival with numerous longitudinal predictors. The International Journal of Biostatistics. DOI: 10.1515/ijb-2025-0049.
  • This document contains the accepted version of the manuscript. The final version can be freely downloaded (Open Access) from the website of the IJB, using this link: https://doi.org/10.1515/ijb-2025-0049.
@@ 1 Introduction
Traditional approaches to the dynamic prediction of survival outcomes include last observation carried forward (LOCF) landmarking (Van Houwelingen, , 2007) and joint modelling (Hickey et al., , 2016). The major advantages of the LOFC approach are its simplicity and ease of implementation; possible disadvantages include the fact that the last observation taken before the landmark may be outdated, and the lack of a model to describe the evolution over time of the longitudinal covariates and to account for measurement error (which may be considerable in the case of biomarkers). Unlike LOCF landmarking, joint models (JMs) enable efficient use of the longitudinal data, and to account for measurement error. However, the estimation of JMs is computationally difficult and demanding, particularly when attempting to include many longitudinal covariates in the JM. To date, high computing times and frequent convergence problems have limited the applicability of JMs to problems with a limited number of longitudinal covariates (Fournier et al., , 2019; Spreafico and Ieva, , 2021), forcing researchers working on data with numerous longitudinal covariates to preselect a subset of those covariates before fitting the JM.
@@
Over the last decades, technological advancements have led to a growing availability of studies where many longitudinal covariates are measured alongside a survival outcome. Examples include observational studies were patients are monitored for several years to study the occurrence of a slowly progressing disease such as cancer (Patel et al., , 2017), dementia (Bennett et al., , 2018) or Alzheimer’s disease (Weiner et al., , 2010), as well as studies that seek to identify biomarkers associated with disease progression and the occurrence of disease milestones among a large number of longitudinal omic variables (Signorelli et al., , 2020; Filbin et al., , 2021). Recently, this increasing availability of repeated measurement data in biomedical studies has driven the development of new statistical methods (Li and Luo, , 2019; Signorelli et al., , 2021; Lin et al., , 2021; Devaux et al., , 2023; Gomon et al., , 2024) and software (Signorelli, , 2024; Devaux et al., , 2024) for dynamic prediction that can better handle the availability of a large number of longitudinal covariates. These methods approach the dynamic prediction problem through a multi-step modelling approach: first, they employ either mixed-effects models (MEMs) or multivariate functional principal component analysis (MFPCA) to describe the evolution over time of the longitudinal covariates and to obtain time-independent summaries of the repeated measurement data; then, they model the relationship between the survival outcome and the available covariates using either Cox models or random survival forests (RSFs).
@@
The goal of this study is to investigate these open questions using publicly available real-world datasets. We gathered data from three longitudinal studies that differ substantially in terms of sample size, number of longitudinal covariates and length of follow-up. We employed four multi-step dynamic prediction methods, alongside two simpler prediction approaches, to predict a relevant time-to-event outcome in each dataset, and proceeded to evaluate the predictive performance of each method according to three different metrics for survival data. Although an additional comparison with the JM approach would be interesting, JMs are not included in this benchmarking study because their estimation through state of the art statistical software for JMs (such as joineMRL and JMbayes2) repeatedly failed when attempting to include more than three to five longitudinal predictors in the JM.
@@ 2.1 Dynamic prediction of survival outcomes
We consider the setup of a longitudinal study where $n$ subjects are enrolled, and interest lies in predicting a time-to-event outcome. Each subject $i\in\{1,2,...,n\}$ is followed from time $t=0$ (study entry / baseline visit) until they either experience the event of interest at time $t_{i}$, or they are right-censored at time $c_{i}$. We denote by $T_{i}$ and $C_{i}$ the random variables corresponding to the event and censoring times, respectively, and let $T_{i}^{*}=\min{\left(T_{i},C_{i}\right)}$. Furthermore, we denote by $t_{i}^{*}$ the observed value of $T_{i}^{*}$, and introduce the censoring indicator $\delta_{i}$ that is $1$ if for subject $i$ the event of interest is observed at $t_{i}^{*}=t_{i}$, and 0 in case of censoring, i.e., if the event has not yet been observed at $t_{i}^{*}=c_{i}$.
@@ 2.3.6 Dynamic Random Survival Forest (DynForest)
2.3.7 Software implementation
All computations presented in this study were performed using R on an AMD EPYC 7662 processor with 2 GHz CPU. After properly subsetting the data as appropriate for the static Cox model and LOCF landmarking model, we proceeded to implement such models in R, relying on the coxph function from R package survival (Therneau and Grambsch, , 2000) for model estimation. We used on the R packages pencal (Signorelli, , 2024) and DynForest (Devaux et al., , 2024) to implement and estimate the PRC and DynForest models. Lastly, to estimate MFPCCox and FunRSF we proceeded to adapt as appropriate the code provided in Li and Luo, (2019) and Lin et al., (2021), respectively. All scripts created to perform the analyses presented in this paper are publicly available at https://github.com/mirkosignorelli/comparisonDynamicPred.
@@ 3.4.2 Performance measures and validation of predictive performance
To evaluate predictive performance, we consider the three most commonly-used measures of predictiveness in survival analysis, namely the Brier score (Graf et al., , 1999), the time-dependent AUC (Heagerty et al., , 2000), and the C index (Pencina and D’Agostino, , 2004).
We employed repeated cross-validation (RCV) to obtain unbiased estimates of the predictive performance of the different models. For the ROSMAP and ADNI datasets we employed 10-fold cross-validation (CV) with 10 repetitions; for the PBC2 dataset, where using 10-fold CV led to very small validation folds, we employed 5-fold CV with 20 repetitions.
@@ 3.4.3 Implementation details
MFPCCox and FunRSF rely on MFPCA to summarize the longitudinal covariates into MFPCA scores. In principle, MFPCA can handle longitudinal data where the measurement times differ across subjects. However, highly-irregular measurement times complicate the estimation and inversion of the univariate covariance functions $C_{q}(t,t^{\prime})$, often leading to numeric errors when attempting to perform the MFPCA step as implemented in MFPCCox and FunRSF. Given that visits in the ADNI and ROSMAP datasets follow a regular measurement grid, but in a few instances the actual visit takes place a few days before or after the planned date, we proceeded to align the measurement times across subjects so as to obtain a regular measurement grid that allowed to solve most of the problems with the estimation of the MFPCA scores. Furthermore, the MFPCA step within MFPCCox and FunSRF requires the specification of the minimum percentage of variance explained (PVE) to retain when reducing the longitudinal covariates to univariate FPCA scores (PVE1), and then from univariate to multivariate FPCA scores (PVE2). We proceeded to set $_{1}= PVE$$_{2}=90\%$ for both methods.
@@ 4.1 Predictive performance on the ADNI data
C index. As shown in Table 1, for all landmark times PRC consistently achieves the highest performance, and FunRSF the lowest one. The relative performance of the remaining methods varies slightly across landmarks: at landmarks 2 and 3, landmarking and static Cox have the second and third highest C index, and are followed by DynForest and MFPCCox. Instead, at landmark 4 DynForest is the second best method. Overall, the relative performance of DynForest and MFPCCox improves as the landmark time increases, whereas that of landmarking and static Cox worsens.
@@
Effect of sample size and number of predictors on predictive performance. To investigate the effect of the sample size $n$ on predictive performance, we proceeded to fit the same models on a third and two thirds of the available observations, considering the first landmark (Supplementary Table 12 and Supplementary Figures 4-5). Moreover, we investigated the effect of the number of predictors $P$ and $Q$ on predictive performance by refitting the same models using a third and two thirds of the covariates considered in the full analysis (Supplementary Table 13 and Supplementary Figures 6-7).
Overall, the sample size does not affect substantially the cross-validated values of the C index, time-dependent AUC and Brier score, however we can notice that the standard deviation over the repetitions of the CV decreases substantially with $n$; this reflects the fact that as the sample size increases, more information is available for accurate evaluation of the predictive performance. On the other hand, increasing the number of predictors leads to higher predictive performance, producing slight increases in the C index and tdAUC, and reductions in the Brier score. This is in line with the expectation that using more predictors can lead to gains in predictive performance (although in this specific dataset the gains appear to be rather small, suggesting that it is already possible to achieve a high predictive performance with a rather small number of predictors).
@@ 4.4 Computing time
We now turn our attention to computing time. Tables 4, 5, and 6 report the average computing time required to estimate each model in k1k-1 folds and to obtain predictions for the remaining fold (averaged over all RCV iterations). Computing time is generally higher for the ROSMAP dataset, where we have the largest sample size and number of predictors, and lowest for the PBC2 dataset where $n$, $P$ and $Q$ are small. By comparing computing time across landmarks, we observe that the computing time typically decreases as the landmark increases; this is simply an artifact of the fact that at later landmarks less individuals are still at risk, reducing the sample size used for model training and prediction.
@@ 5.3 Software and computing time
Supplementary material
Supplementary tables and figures are available in the “Supplementary Material” pdf file.
Code and data
The R code used to perform the analyses described in this paper is available at
https://github.com/mirkosignorelli/comparisonDynamicPred. A copy of the PBC2 data is available at the same URL.
The data from the ROSMAP study are available upon motivated request from the Rush Alzheimer’s Disease Center (RADC) of the RUSH University. ROSMAP resources can be requested at https://www.radc.rush.edu and www.synpase.org.
The ADNI data are available upon motivated request from the Alzheimer’s Disease Neuroimaging Initiative (ADNI) database (http://adni.loni.usc.edu).
@@ Acknowledgements
ROSMAP. Data from the ROSMAP study were obtained from the Rush Alzheimer’s Disease Center (RADC) of the RUSH University (https://www.radc.rush.edu). ROSMAP is supported by P30AG10161, P30AG72975, R01AG15819, R01AG17917, U01AG46152, and U01AG61356.
ADNI. Data from the ADNI study were obtained from the ADNI database (http://adni.loni.usc.edu). As such, the investigators within the ADNI contributed to the design and implementation of ADNI and/or provided data but did not participate in analysis or writing of this paper. A complete listing of ADNI investigators can be found at https://adni.loni.usc.edu/wp-content/uploads/how_to_apply/ADNI_Acknowledgement_List.pdf.